Two truckers leave San Francisco at the same time. They take 10 and 28 days, respectively, to reach their destination and return to San Francisco. The truckers each take continuous trips to and from San Francisco. How many days will pass before the two truckers leave San Francisco on the same day again?
Found 3 solutions by ikleyn, ankor@dixie-net.com, saw:
Answer by ikleyn(53763)

.

The Least Common Multiple (LCM) of the numbers 10 and 28 is 2^2*5*7 = 140.


ANSWER.  140 days.

Solved.


Answer by ankor@dixie-net.com(22740)

Two truckers leave San Francisco at the same time.
They take 10 and 28 days, respectively, to reach their destination and return to San Francisco.
The truckers each take continuous trips to and from San Francisco.
How many days will pass before the two truckers leave San Francisco on the same day again?
:
find the least common multiple of 10 and 28
using prime factors
10: 2,5
28: 2,2,7
least common multple
2 * 2 * 5 * 7 = 140 days

Answer by saw(34)

Solution.
Find the Least Common Multiple by using prime factors
10 = 2*5
28 = 2*2*7
lcm(10,28) = 2*2*5*7 = 140
So the answer is 140 days.